# Practical indistinguishability in a gene regulatory network inference problem, a case study

### v1

@@ 3 Results
To infer structural features of the regulatory network that underpins the developmental decision in a wild-type genetic background, we fit 13,824 different model structures to RNA-seq data from the wild-type experiment. A subset of around 1000 models gives a reasonable fit to data (Figure 4 A). After about 1000 models, the cost function curve over model structures plateaus, before increasing dramatically (see Figure 4 A). The best-fitting models give a reasonable description of the data, in that the estimate closely describes the data and the uncertainty is more constrained than other model structures in the family we considered (Figure 4B, bottom). The model structures that result in the highest cost function values do not describe the data well and are highly uncertain (Figure 4B, top). In particular, the estimate for nhr-40 mRNA is almost entirely unconstrained. Qualitatively, the cost function curve seems to have three distinct regions: models that describe the gene expression data reasonable well (pink in Figure 4 A), models that describe the gene expression data more poorly but are not dominated by uncertainty (gray before the dashed line in Figure 4 A), and models that do not capture the data and are dominated by uncertainty (gray after the dashed line in Figure 4 A). As expected, there is a linear relationship between the model’s cost-function-value and the estimate’s uncertainty (Figure 4C).
@@
To learn which characteristics distinguish better fitting models, we used a decision tree to classify models that produced an acceptable fit (Figure 4 A, pink) from those that produced an unacceptable fit (Figure 4 A, gray), using a broad array of 21 distinct structural features of the models (see Section 6.6). We find that the models that generate the best, more certain fits to data are enriched in several model features (Figure 4 D & E, pink). In particular, NHR-40 likely regulates eud-1 and nhr-40 and sometimes sult-1. We also found that eud-1 and nhr-40 are likely positively regulated, and nhr-40 is likely autoregulated. We found that 325 regulated networks that share these features all generate an acceptable fit to the data (example in Figure 4D, right).
@@
Notably, not all models that generate an acceptable fit to the data have these characteristics. Roughly 35% of the model structures that produce an acceptable fit share these features. The analysis also revealed a broad array of 21 model structures that predictably generate unacceptable fits to the data (Figure 4 E, gray). For many models, the decision tree could not predict whether a model would yield an acceptable or unacceptable fit. This is expected because the plateau of the cost function value curve (Figure 4 A) is relatively rounded, so models on either side of the threshold we selected likely share common features. A handful of model structures resulted in many of the misclassifications (Figure 4D, dotted gray). These structures are closely related to network structures that produce both acceptable and unacceptable fits to the data.
@@
Figure 4: A: We fit 13,824 models corresponding to different gene regulatory network structures to WT eud-1, nhr-40, sult-1 gene expression. Fitting each mathematical model to the data results in a cost function value. We sort the models by their cost function value. The curve that results has a noticeable plateau after about 1000 models (pink). B: Forward simulations from the discovered parameters for a well-fit (bottom) and a poor-fitting model (top). The dynamics of eud-1 mRNA (purple), nhr-40 mRNA (orange), sult-1 mRNA (blue). The uncertainty (one standard deviation) is shown as the band of the corresponding color. C: Estimated standard deviation of the noise vs. the cost function value. As expected, poorer fits have a higher estimated noise. D: Left, legend of symbols used. Middle: an example of a network class discovered in our analysis. This network class has 189 distinct regulatory networks within which each shares positive regulation of eud-1 and nhr-40, regulation of eud-1 by NHR-40, autoregulation of nhr-40, and regulation of sult-1 by NHR-40. Right: one of the 189 networks that occupy the network class. This network is the unified regulatory network we describe in Figure 5. E: Discovery of the model set. We used a decision tree to classify acceptable fits (pink in A) from unacceptable fits (gray in A). We discovered network features shared by models that generate acceptable fits (pink) and unacceptable fits (gray). We also find several combinations of model features that do not neatly predict acceptable or unacceptable fits. This is likely due to the rounded nature of the plateau in A. Note that only the features used in classification are shown in this diagram.
@@
When we fit a wide range of model structures to the eud-1 KO experiment, we see that only a handful of models achieve a good fit (see Figure 5A, middle), and the cost function curve exhibits a large gap to models with poorer fits. The two models with the lowest cost function values have very similar model structures. In particular, nhr-40 is positively regulated by NHR-40 and sult-1 is positively regulated by NHR-40 and SULT-1. The difference between the models is whether or not SULT-1 positively regulates nhr-40. When we fit a wide range of model structures to sult-1 KO experiment, we notice that about 10 models seem to give a reasonable fit to the data before the cost function curve plateaus (see Figure 5A, right). These models tend to include positive regulation of eud-1 by NHR-40, autoregulation of eud-1 (positive or negative), and either positive autoregulation of nhr-40 or positive regulation of nhr-40 by EUD-1. We then probe the model ensembles associated with each of the experiments (Figure 5 A, pink) for a common regulatory network structure that can explain all three experiments. We find a model structure (Figure 5 A, orange dot and Figure 5 C) with strong commonalities to the model we did synthetic testing on. Both models feature a high degree of positive regulation of the network and include regulation of eud-1 by NHR-40 and autoregulation of nhr-40. The unified model structure has the form
[eud-1 mRNA], Observed: $\dfrac{dM_{e}}{dt}=k_{t_{e_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{1}}+\hat{E}}\right)\left(\dfrac{\hat{N}}{\hat{K_{4}}+\hat{N}}\right)-k_{-m}M_{e},$ $M_{e}(0)=M_{eo}$ (13)
[scaled EUD-1 protein], Unobserved: $\dfrac{d\hat{E}}{dt}=M_{e}-k_{-p}\hat{E}$, $\hat{E}(0)=\hat{E_{o}}$ (14)
[nhr-40 mRNA], Observed: $\dfrac{dM_{n}}{dt}=k_{t_{n_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{2}}+\hat{E}}\right)\left(\dfrac{\hat{N}}{\hat{K_{5}}+\hat{N}}\right)-k_{-m}M_{n},$ $M_{n}(0)=M_{no}$ (15)
[scaled NHR-40 protein], Unobserved: $\dfrac{d\hat{N}}{dt}=M_{n}-k_{-p}\hat{N}$, $\hat{N}(0)=\hat{N_{o}}$ (16)
[sult-1 mRNA], Observed: $\dfrac{dM_{s}}{dt}=k_{t_{s_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{3}}+\hat{N}}\right)\left(\dfrac{\hat{S}}{\hat{K_{6}}+\hat{S}}\right)-k_{-m}M_{s},$ $M_{s}(0)=M_{so}$ (17)
[scaled SULT-1 protein], Unobserved: $\dfrac{d\hat{S}}{dt}=M_{s}-k_{-p}\hat{S}$, $\hat{S}(0)=\hat{S_{o}}$, (18)
@@
where $M_{e}$, $\hat{E}$, $M_{n}$, $\hat{N}$, $M_{s}$, and $\hat{S}$ is the concentration of eud-1 mRNA, the scaled concentration of EUD-1 protein, the concentration of nhr-40 mRNA, the scaled concentration of NHR-40 protein, the concentration of sult-1 mRNA, and the scaled concentration of SULT-1 protein, respectively. As predicted by the decision tree, the model is dominated by positive regulation and features regulation of eud-1 by NHR-40 and autoregulation of nhr-40. When sult-1 is knocked out, the model reduces to Eqs. (13)-(16). When eud-1 is knocked out, the model reduces to the following equations,
[eud-1 mRNA], Observed: $\dfrac{dM_{e}}{dt}=k_{t_{e_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{1}}+\hat{E}}\right)\left(\dfrac{\hat{N}}{\hat{K_{4}}+\hat{N}}\right)-k_{-m}M_{e},$ $M_{e}(0)=M_{eo}$ (19)
[scaled EUD-1 protein], Unobserved: $\dfrac{d\hat{E}}{dt}=M_{e}-k_{-p}\hat{E}$, $\hat{E}(0)=\hat{E_{o}}$ (20)
[nhr-40 mRNA], Observed: $\dfrac{dM_{n}}{dt}=k_{t_{n_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{5}}+\hat{N}}\right)-k_{-m}M_{n},$ $M_{n}(0)=M_{no}$ (21)
[scaled NHR-40 protein], Unobserved: $\dfrac{d\hat{N}}{dt}=M_{n}-k_{-p}\hat{N}$, $\hat{N}(0)=\hat{N_{o}}$. (22)
@@
Here, we attempted to infer the structure of a regulatory network that controls the dimorphic mouth form decision in P. pacficus using RNA-seq data from three experiments. We fit a total of 13,824 models to the wild type experiment, and 144 models to each of the knock-out experiments, and found a model set, or combinations of features that are shared across the models that generate acceptable fits to the data, for each experiment. Collating these model sets, we put forth a unified regulatory network structure that exists in each of the model sets associated with the experiment (see Figure 5). The regulatory network model is relatively complex and features a high degree of positive autoregulation and positive cross-regulation. It is important to note that this is one possible regulatory motif that explains the data reasonably well across experiments, but a very high degree of uncertainty still exists given the size of the model sets associated with the wild type and sult-1 knock out experiment—roughly 1000 models and 10 models, respectively. More experimental work is needed to refine our model clouds and verify the possible regulatory network structure that we identified in our analysis.
Figure 5: A: Cost function curves sorted by value for the WT experiment, the eud-1 knock-out experiment, and the sult-1 knock-out experiment. Pink shaded regions indicate the best fits up to a plateau in the curve. B: Commonality in the regulatory network structure for the best fitting models (pink shaded region in A). C: An example unified regulatory network. The model structure appears in (A) in orange. D: Model simulations of the unified regulatory mechanism across experiments at optimized parameter values (only one replicate shown). Scaled eud-1 mRNA, nhr-40 mRNA, sult-1 mRNA expression is shown as purple, orange, and blue dots, respectively. Model simulations at optimization parameter values are shown as a solid line in the corresponding color. The uncertainty (two standard deviations) is shown as the band of the corresponding color.
@@ 4 Discussion
Here we performed a case study of practical indistinguishability across 13,824 distinct model structures to infer features of a regulatory network that underpins a developmental decision in the nematode P. pacificus. We found model sets, or combinations of features that are shared across the models that generate acceptable fits to the data, for three experiments. We then found a regulatory network that is shared across each of the model sets and is capable of describing the data across experimental conditions. This regulatory network exhibits a high degree of positive regulation and interconnectivity between the key regulators, eud-1, sult-1, and nhr-40. Mathematically, this analysis is a natural complement to emerging theoretical results in structural indistinguishability. There are several natural next steps to be addressed on the biological and mathematical side.
@@
To further refine our understanding of the underlying mechanism of the developmental decision based on what we learned in this in silico analysis, there are several next steps to be addressed in future work. Across model sets, we found that NHR-40 was deeply involved in the regulation of the effective networks we discovered. A deeper understanding of the regulatory network could be found through an nhr-40 knock-out experiment, to understand the dynamics of the system in the absence of this key regulator. Additionally, one could use the discovered model sets to perform optimal experiment design to suggest new experiments that limit cost and maximize our ability to distinguish the models associated with each experiment. Possible informative experiments include more frequently sampled gene expression data, measurements of protein states, or applying other perturbations to system states that have an indirect effect on the genes of interest.
@@
The analysis we presented here examined practical indistinguishability across non-linear model structures using experimental data. Mathematically, it would be of interest to extend the structural indistinguishability results to nonlinear systems and understand the connections between practical identifiability and practical distinguishability. Structural indistinguishability is a data-independent approach to understanding which models can be distinguished from one another using the input-output equation or graph theoretic approaches [21, 22]. These theoretical results are largely limited to linear models and cannot currently be applied to non-linear systems, such as Eqs. (13)-(14) and do not consider the impact of experimental data on indistinguishability. A natural next step is to attempt to extend the theoretical results for linear systems to systems with non-linear forms, such as Eqs. (13)-(14). Practical identifiability examines if unique parameters can be found given the associated data. An additional future direction would be to extend frameworks for practical identifiability to examine and identify sources of structural uncertainty in the model, or identify additional terms that would retain practical identifiability given the data.
@@
We also assume that the broad dynamics of the regulatory network are deterministic and can be described by a deterministic model. However, it is known that gene expression can be stochastic [29]. Our efforts are a first pass at beginning to understand the underlying regulatory network, and inference of a stochastic system would not be possible without more sample time points and likely single-cell measurements. Given the highly conserved nature of this developmental pathway and the fact that we are analyzing bulk RNA-seq rather than data from single cells, it is perhaps reasonable to expect a relatively deterministic genetic signal at the organismal level.
@@
were estimated between 1x10-9 and 1x108+1. Narrower parameter spaces tended to lead to poor fits or fits in which parameters often hit the bounds during optimization. Clearly, there is a trade-off between the size of the parameter space and our ability to efficiently search it using multi-start optimization. The fact that models with similar structures provide the best fits to the data provides some reassurance that we can navigate this space in a reasonable manner. In our analysis, we noticed that many of the best fits across model structures included a relatively high estimate for $\hat{N}(0)$, the initial condition for the scaled NHR-40 protein, on the order of magnitude of 1x107. The only way that $\hat{N}(0)$ could be on the order of magnitude of 1x107 is if $\hat{N}(0)$ is relatively large and the translation rate is relatively small. It is feasible and perhaps even likely that additional biological processes not captured in our model are lowering the effective translation rate, such as a post-transcriptional modification or RNA degradation that is beyond the scope of the model and current understanding of the system.
@@
In the model sets associated with each of the experiments, NHR-40 was implicated as a regulator of the network, but this may be an artifact of key modeling assumptions we were forced to make due to the constraints in the data. Based on the experimental data, nhr-40 gene expression was much higher than either eud-1 or sult-1 in all experiments. In an attempt to balance the biological reality and limit the number of parameters we attempt to estimate, we did not attempt to estimate the initial condition for EUD-1 or SULT-1. We assume that if gene expression is relatively low, as it is for eud-1 or sult-1 across all experiments, then the associated protein concentration should also be relatively low. Therefore, we fixed the EUD-1 and SULT-1 protein initial condition to a low pre-set value during parameter estimation (see Methods). Our structural identifiability analysis suggested that the initial condition for the scaled NHR-40 protein is identifiable if it appeared in at least one of the mRNA production terms. For model structures that included NHR-40 as a regulator of the network, we sought to estimate the initial condition, $\hat{N}(0)$. It may not be surprising that models that best fit the data often included NHR-40 somewhere in the regulatory network, as these models included one additional parameter, $\hat{N}(0)$. However, we inferred more features of the regulatory network than just the presence of NHR-40 as a regulator in our analysis (see Figure 5).
@@
This work was relatively computationally heavy. After trial and error, we ran parameter estimation using 1000 seeds for each model. It is possible that the results could slightly change if we were able to re-run our analysis using 10,000, 100,000, or more seeds, though we obtained similar results when we spot-checked a handful of models using 10,000 seeds. Increasing to 10,000, 100,000, or more seeds for all models was infeasible given our current computing resources and the number of model structures we wanted to probe in depth.
@@
An additional caveat relates to our choice of error model. The biological data are relatively noisy across replicates, especially when gene expression is higher. The number of replicates and time points - while standard, if not better than many biological data sets - did not enable us to determine the error model for the system in a data-driven manner, and there was no clear first principles approach to determining the error model using knowledge of the system. The data may also include some error in time or the “x-direction,” which can occur if a replicate is slightly ahead or behind others, and further complicates the error model. We did not attempt to correct for this asynchrony. Ultimately, we assumed a multiplicative error model, largely based on qualitative features of the data, and optimized a maximum likelihood estimate. We found similar qualitative modeling results when we optimized a weighted least squares cost function and prescribed the weight to be inversely related to the variance across replicates.
@@
Here, we attempted to probe the limits of mechanistic inference using data-driven modeling applied to biological data in a common yet challenging data regime. We found that we could partially resolve features of the effective regulatory network that controls the dimorphic mouth-form decision in P. pacficus, but could not completely discern the structure of the regulatory network given the limitations of the data, our current understanding of the underlying biology, and our computational constraints. Our data-driven modeling results identify features of the regulatory network that generate acceptable fits to the known data and suggest future experiments, such as temporally sampling expression of eud-1 and sult-1 in a loss of function nhr-40 knockout over the course of development or using the discovered model sets to perform optimal experimental design to suggest sampling times, to further refine features of the regulatory network. In turn, our analysis can be iteratively run as new data becomes available, potentially narrowing the size of our model sets and improving our ability to distinguish model structures. Beyond the biological implications of our work, this analysis attempts to understand the sensitivity of our results as a function of model structure and probes what can be mechanistically inferred based on typical biological data. These analyses can be applied to any system where structural uncertainty poses a challenge.

### v2

@@ 3 Results
To infer structural features of the regulatory network that underpins the developmental decision in a wild-type genetic background, we fit 13,824 different model structures to RNA-seq data from the wild-type experiment. A subset of around 1000 models gives a reasonable fit to data (Figure 4 A). After about 1000 models, the cost function curve over model structures plateaus, before increasing dramatically (see Figure 4 A). The best-fitting models give a reasonable description of the data, in that the estimate closely describes the data and the uncertainty is more constrained than other model structures in the family we considered (Figure 4B, bottom). The model structures that result in the highest cost function values do not describe the data well and are highly uncertain (Figure 4B, top). In particular, the estimate for nhr-40 mRNA is almost entirely unconstrained. Qualitatively, the cost function curve seems to have three distinct regions: models that describe the gene expression data reasonably well (pink in Figure 4 A), models that describe the gene expression data more poorly but are not dominated by uncertainty (gray before the dashed line in Figure 4 A), and models that do not capture the data and the uncertainty is as large as the dynamic range of the data (gray after the dashed line in Figure 4 A). As expected, there is a positive correlation between the model’s cost-function-value and the estimate’s uncertainty (Figure 4C).
@@
To learn which characteristics distinguish better fitting models, we used a decision tree to classify models that produced an acceptable fit (Figure 4 A, pink) from those that produced an unacceptable fit (Figure 4 A, gray), using a broad array of 21 distinct structural features of the models (see Section 6.6). We find that the models that generate the best, more certain fits to data are enriched in several model features (Figure 4 D, E, F ). In particular, NHR-40 likely regulates eud-1 and nhr-40 and sometimes sult-1. We also found that eud-1 and nhr-40 are likely positively regulated, and nhr-40 is likely autoregulated. We found that 325 regulated networks across two classes that all generate an acceptable fit to the data (Figure 4D).
@@
Notably, not all models that generate an acceptable fit to the data have these characteristics. Roughly 35% of the model structures that produce an acceptable fit share these features (see Figure 6). The analysis also revealed a broad array of 21 model structures that predictably generate unacceptable fits to the data (see Figure 6). For some models, the decision tree could not predict whether a model would yield an acceptable or unacceptable fit (see Figure 6). This is expected because the plateau of the cost function value curve (Figure 3 A) is relatively rounded, so models on either side of the threshold we selected likely share common features. A handful of model structures resulted in many of the misclassifications (see Figure 6). These structures are closely related to network structures that produce both acceptable and unacceptable fits to the data.
@@
Figure 4: A: We fit 13,824 models corresponding to different gene regulatory network structures to WT eud-1, nhr-40, sult-1 gene expression. Fitting each mathematical model to the data results in a cost function value. We sort the models by their cost function value. The curve that results has a noticeable plateau after about 1000 models (pink). B: Forward simulations from the discovered parameters for a well-fit (bottom) and a poor-fitting model (top). The dynamics of eud-1 mRNA (purple), nhr-40 mRNA (orange), sult-1 mRNA (blue). The uncertainty (one standard deviation) is shown as the band of the corresponding color. C: Estimated standard deviation of the noise vs. the cost function value. As expected, poorer fits have a higher estimated noise. D: We use a decision tree classify acceptable fits (pink in A) and unacceptable fits (gray in A) based on structural features of the 13,824 models. We find two classes of model structures that predict acceptable fits. The first class includes positive regulation of eud-1 and nhr-40, regulation of eud-1 by NHR-40, autoregulation of nhr-40 and no regulation of sult-1 by NHR-40. The second class (pink dashed box) includes positive regulation of eud-1 and nhr-40, regulation of eud-1 by NHR-40, autoregulation of nhr-40 and regulation of sult-1 by NHR-40. E: Resolved features of the second class (dashed pink square) of the model set (D) are shown in black and the unresolved features as shown as dashed brown arrows. F: An specific example regulatory network structure from the example class shown in E. This network features a high degree of positive autoregulation and positive cross-regulation.
@@
When we fit a wide range of model structures to the eud-1 KO experiment, we see that only a handful of models achieve a good fit (see Figure 5A, middle), and the cost function curve exhibits a large gap to models with poorer fits. The two models with the lowest cost function values have very similar model structures. In particular, nhr-40 is positively autoregulated and sult-1 is positively regulated by NHR-40 and positively autoregulated. The difference between the models is whether or not SULT-1 positively regulates nhr-40. When we fit a wide range of model structures to sult-1 KO experiment, we notice that about 10 models seem to give a reasonable fit to the data before the cost function curve plateaus (see Figure 5A, right). These models tend to include positive regulation of eud-1 by NHR-40, autoregulation of eud-1 (positive or negative), and either positive autoregulation of nhr-40 or positive regulation of nhr-40 by EUD-1. We then probe the model sets associated with each of the experiments (Figure 5 A, pink) for a common regulatory network structure that can explain all three experiments. We find a model structure (Figure 5 A, orange dot and Figure 5 C) with strong commonalities to the model we did synthetic testing on. Both models feature a high degree of positive regulation of the network and include regulation of eud-1 by NHR-40 and autoregulation of nhr-40. The unified model structure has the form
[eud-1 mRNA], Obs.: $\dfrac{dM_{e}}{dt}=\alpha_{1}+k_{t_{e_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{1}}+\hat{E}}\right)\left(\dfrac{\hat{N}}{\hat{K_{4}}+\hat{N}}\right)-k_{-m}M_{e},$ $M_{e}(0)=M_{eo}$ (13)
[scaled EUD-1 protein], Unobs.: $\dfrac{d\hat{E}}{dt}=M_{e}-k_{-p}\hat{E}$, $\hat{E}(0)=\hat{E_{o}}$ (14)
[nhr-40 mRNA], Obs.: $\dfrac{dM_{n}}{dt}=\alpha_{2}+k_{t_{n_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{2}}+\hat{E}}\right)\left(\dfrac{\hat{N}}{\hat{K_{5}}+\hat{N}}\right)-k_{-m}M_{n},$ $M_{n}(0)=M_{no}$ (15)
[scaled NHR-40 protein], Unobs.: $\dfrac{d\hat{N}}{dt}=M_{n}-k_{-p}\hat{N}$, $\hat{N}(0)=\hat{N_{o}}$ (16)
[sult-1 mRNA], Obs.: $\dfrac{dM_{s}}{dt}=\alpha_{3}+k_{t_{s_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{3}}+\hat{N}}\right)\left(\dfrac{\hat{S}}{\hat{K_{6}}+\hat{S}}\right)-k_{-m}M_{s},$ $M_{s}(0)=M_{so}$ (17)
[scaled SULT-1 protein], Unobs.: $\dfrac{d\hat{S}}{dt}=M_{s}-k_{-p}\hat{S}$, $\hat{S}(0)=\hat{S_{o}}$, (18)
@@
where $M_{e}$, $\hat{E}$, $M_{n}$, $\hat{N}$, $M_{s}$, and $\hat{S}$ is the concentration of eud-1 mRNA, the scaled concentration of EUD-1 protein, the concentration of nhr-40 mRNA, the scaled concentration of NHR-40 protein, the concentration of sult-1 mRNA, and the scaled concentration of SULT-1 protein, respectively. As predicted by the feature importance analysis using the decision tree decision tree, the model is dominated by positive regulation and features regulation of eud-1 by NHR-40 and autoregulation of nhr-40. When sult-1 is knocked out, the model reduces to Eqs. (13)-(16). When eud-1 is knocked out, the model reduces to the following equations,
[eud-1 mRNA], Obs.: $\dfrac{dM_{e}}{dt}=\alpha_{1}+k_{t_{e_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{1}}+\hat{E}}\right)\left(\dfrac{\hat{N}}{\hat{K_{4}}+\hat{N}}\right)-k_{-m}M_{e},$ $M_{e}(0)=M_{eo}$ (19)
[scaled EUD-1 protein], Unobs.: $\dfrac{d\hat{E}}{dt}=M_{e}-k_{-p}\hat{E}$, $\hat{E}(0)=\hat{E_{o}}$ (20)
[nhr-40 mRNA], Obs.: $\dfrac{dM_{n}}{dt}=\alpha_{2}+k_{t_{n_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{5}}+\hat{N}}\right)-k_{-m}M_{n},$ $M_{n}(0)=M_{no}$ (21)
[scaled NHR-40 protein], Unobs.: $\dfrac{d\hat{N}}{dt}=M_{n}-k_{-p}\hat{N}$, $\hat{N}(0)=\hat{N_{o}}$. (22)
@@
Here, we attempted to infer the structure of a regulatory network that controls the dimorphic mouth form decision in P. pacficus using RNA-seq data from three experiments. We fit a total of 13,824 models to the wild type experiment, and 144 models to each of the knock-out experiments, and found a model set, or combinations of features that are shared across the models that generate acceptable fits to the data, for each experiment (see Figure 5B). Collating these model sets, we put forth a unified regulatory network structure that exists in each of the model sets associated with the experiment (see Figure 5). The regulatory network model is relatively complex and features a high degree of positive autoregulation and positive cross-regulation (see Figure 5C). It is important to note that this is one possible regulatory motif that explains the data reasonably well across experiments, but a very high degree of uncertainty still exists given the size of the model sets associated with the wild type and sult-1 knock out experiment—roughly 1000 models and 10 models, respectively. More experimental work is needed to refine our model sets and verify or exclude the possible regulatory network structure that we identified in our analysis.
Figure 5: A: Cost function curves sorted by value for the WT experiment, the eud-1 knock-out experiment, and the sult-1 knock-out experiment. Pink shaded regions indicate the best fits up to a plateau in the curve. B: Commonality in the regulatory network structure for the best fitting models (pink shaded region in A). C: An example unified regulatory network. The model structure appears in (A) in orange. D: Model simulations of the unified regulatory mechanism across experiments at optimized parameter values (only one replicate shown). Scaled eud-1 mRNA, nhr-40 mRNA, sult-1 mRNA expression is shown as purple, orange, and blue dots, respectively. Model simulations at optimization parameter values are shown as a solid line in the corresponding color. The uncertainty (one standard deviations) is shown as the band of the corresponding color.
@@ 4 Discussion
Here we performed a case study of practical indistinguishability across 13,824 distinct model structures to infer features of a regulatory network that underpins a developmental decision in the nematode P. pacificus. We found model sets, or collections of models with shared regulatory network features that best fit the data, for each of the three experiments (see Figures 4 and 5). We then found a regulatory network in the intersection of the three model sets and is capable of describing the data across each of the experimental conditions (see Figure 5C and D). This regulatory network exhibits a high degree of positive regulation and interconnectivity between the key regulators, eud-1, sult-1, and nhr-40. There are several natural next steps to be addressed on the biological and mathematical side.
@@
To further refine our understanding of the underlying mechanism of the developmental decision based on what we learned in this in silico analysis, there are several next steps to be addressed in future work. Across model sets, we found that NHR-40 was heavily involved in the regulation of the effective networks we discovered. A deeper understanding of the regulatory network could be found through an nhr-40 knock-out experiment, to understand the dynamics of the system in the absence of this key regulator. Additionally, one could use the discovered model sets to perform optimal experiment design to suggest new experiments that limit cost and maximize our ability to distinguish the models associated with each experiment. Possible informative experiments include more frequently sampled gene expression data, measurements of protein states, or applying other perturbations to system states that have an indirect effect on the genes of interest.
@@
Mathematically, this analysis is a natural complement to emerging theoretical results in structural indistinguishability [30, 31]. The analysis we presented here examined practical indistinguishability across non-linear model structures using experimental data. Mathematically, it would be of interest to extend the structural indistinguishability results to nonlinear systems and understand the connections between practical identifiability, or the ability to find unique parameters given the associated data, and practical indistinguishability. Structural indistinguishability is a data-independent approach to understanding which models can be distinguished from one another using the input-output equation or graph theoretic approaches [30, 31]. These theoretical results are largely limited to linear models and cannot currently be applied to non-linear systems, such as Eqs. (13)-(14) and do not consider the impact of experimental data on indistinguishability. A natural next step is to attempt to extend the theoretical results for linear systems to systems with non-linear forms, such as Eqs. (13)-(14). Practical identifiability examines if unique parameters can be found given the associated data. An additional future direction would be to extend frameworks for practical identifiability to examine and identify sources of structural uncertainty in the model, or identify additional terms that would retain practical identifiability given the data.
@@
We also assume that the broad dynamics of the regulatory network are deterministic and can be described by a deterministic model. However, it is known that gene expression can be stochastic [38]. Our efforts are a first pass at developing a model for the underlying regulatory network, and inference of a stochastic system would not be possible without more sample time points and likely single-cell measurements. Given the highly conserved nature of this developmental pathway and the fact that we are analyzing bulk RNA-seq rather than data from single cells, it is perhaps reasonable to expect a relatively deterministic genetic signal at the organismal level.
@@
were estimated between 1x10-9 and 1x108+1. Narrower parameter spaces tended to lead to poor fits or fits in which parameters often hit the bounds during optimization. There is a trade-off between the size of the parameter space and our ability to efficiently search it using multi-start optimization (see Section 6.5). The fact that models with similar structures provide the best fits to the data provides some reassurance that we can navigate this space in a reasonable manner. In our analysis, we noticed that many of the best fits across model structures included a relatively high estimate for $\hat{N}(0)$, the initial condition for the scaled NHR-40 protein, on the order of magnitude of 1x107. The only way that $\hat{N}(0)$ could be on the order of magnitude of 1x107 is if $\hat{N}(0)$ is relatively large and the translation rate, $k_{t}$, is relatively small. It is feasible and perhaps even likely that additional biological processes not captured in our model are lowering the effective translation rate, such as a post-transcriptional modification or RNA degradation that is beyond the scope of the model and current understanding of the system.
@@
In the model sets associated with each of the experiments, NHR-40 was implicated as a regulator of the network, but this may be an artifact of key modeling assumptions we were forced to make due to the constraints in the data. Based on the experimental data, nhr-40 gene expression was much higher than either eud-1 or sult-1 in all experiments. In an attempt to balance the biological reality and limit the number of parameters we attempt to estimate, we did not attempt to estimate the initial condition for EUD-1 or SULT-1. We assume that if gene expression is relatively low, as it is for eud-1 or sult-1 across all experiments, then the associated protein concentration should also be relatively low. Therefore, we fixed the EUD-1 and SULT-1 protein initial condition to a low pre-set value during parameter estimation (see Methods). Our structural identifiability analysis suggested that the initial condition for the scaled NHR-40 protein is identifiable if it appeared in at least one of the mRNA production terms. For model structures that included NHR-40 as a regulator of the network, we sought to estimate the initial condition, $\hat{N}(0)$. It may not be surprising that models that best fit the data often included NHR-40 somewhere in the regulatory network, as these models included one additional parameter, $\hat{N}(0)$. However, we inferred more features of the regulatory network than just the presence of NHR-40 as a regulator in our analysis (see Figures 4-5).
@@
This work was relatively computationally heavy. After trial and error, we ran parameter estimation using 1000 seeds for each model. It is possible that the results could slightly change if we were able to re-run our analysis using 10,000, 100,000, or more seeds, though we obtained similar results when we spot-checked a handful of models using 10,000 seeds. Increasing to 100,000, or more seeds for all models was infeasible given our current computing resources and the number of model structures we wanted to probe.
@@
An additional caveat relates to our choice of error model. The biological data are relatively noisy across replicates, especially when gene expression is higher. The number of replicates and time points - while standard, if not better than many biological data sets - did not enable us to determine the error model for the system in a data-driven manner, and there was no clear first principles approach to determining the error model using knowledge of the system. The data may also include some error in time or the “x-direction,” which can occur if a replicate is slightly ahead or behind others, and further complicates the error model. We did not attempt to correct for this possible temporal asynchrony. Ultimately, we assumed a multiplicative error model, largely based on qualitative features of the data, and optimized a maximum likelihood estimate (see Section 6.5), though we also found similar qualitative modeling results when we optimized a weighted least squares cost function and prescribed the weight to be inversely related to the variance across replicates.
@@
Here, we attempted to probe the limits of mechanistic inference using data-driven modeling applied to biological data in a common yet challenging data regime. We found that we could partially resolve features of the effective regulatory network that controls the dimorphic mouth-form decision in P. pacficus, but could not completely discern the structure of the regulatory network given the limitations of the data, our current understanding of the underlying biology, and our computational constraints. Our data-driven modeling results identify features of the regulatory network that generate acceptable fits to the known data and suggest future experiments, such as temporally sampling expression of eud-1 and sult-1 in a loss of function nhr-40 knockout over the course of development or using the discovered model sets to perform optimal experimental design to suggest sampling times, to further refine features of the regulatory network. In turn, our analysis can be iteratively run as new data becomes available, potentially narrowing the size of our model sets and improving our ability to distinguish model structures. Beyond the biological implications of our work, this analysis attempts to understand the sensitivity of our results as a function of model structure and probes what can be mechanistically inferred based on typical biological data. These analyses can be applied to any system where structural uncertainty poses a challenge and the experimental data makes algorithmic model selection infeasible.